ROVIBRATIONAL LOWEST BENDING MODE TRANSITIONS OF THE PROPYNYLIDYNE RADICAL

M. CARIS, \underline{T.F. GIESEN}, H.S.P. MÜLLER, F. LEWEN, S. SCHLEMMER, I. PHYSIKALISCHES INSTITUT, UNIVERSITY OF COLOGNE, ZÜLPICHER STR. 77, 50937 KÖLN, GERMANY; C. DUAN, COLLEGE OF PHYSICAL SCIENCE AND TECHNOLOGY, HUAZHONG NORMAL UNIVERSITY, WUHAN 430079, P.R. CHINA; K.M.T. YAMADA, AIST, TSUKUBA-WEST, 16-1 ONOGAWA, TSUKUBA, 305-8569 JAPAN.

Precise measurements on rovibrational transitions of the ν_{4}(^{2}Σ^{μ}) CCH bending mode of (X^{2}Π) C_{3}H have been performed using the Cologne Supersonic Jet Spectrometer for Terahertz Applications (SuJeSTA). The molecules were formed in a supersonic jet expansion when a gas mixture of C_{2}H_{2} and CO diluted in He was introduced to an electrical discharge nozzle. Radiation of a backward wave oscillator (BWO) probes the molecular beam 40 mm downstream and spectra in the frequency range 360-600 GHz were recorded. A strong Renner-Teller coupling effect shifts the (^{2}Σ^{μ}) component of the ν_{4} bending mode towards lower energies and into the submillimeter region. Pure rotational transitions in the ^{2}Π_{1/2} and ^{2}Π_{3/2} ground state levels and in the v_{4} = 1 (^{2}Σ^{μ}) excited vibration state were recorded, extending earlier data published by Yamamoto et al. ^{a} In addition, some rovibrational transitions between the ground state and the v_{4} = 1 (^{2}Σ^{μ}) state have been identified. Thus the vibrational energy of the ν_{4}(^{2}Σ^{μ}) state was determined precisely as 609977.1(42) MHz or 20.34664(14) cm^{−1}. Furthermore, the Coriolis coupling constant between the ground and ν_{4}(^{2}Σ^{μ}) state has also been determined with greatly improved precision.

aS. Yamamoto, S. Saito, H. Suzuki, S. Deguchi, N. Kaifu, and S. Ishikawa, Ap. J. 348, 363-369 (1990)